# Lyapunov exponents and geodesic stability of Schwarzschild black hole in the non-commutative gauge theory of gravity

### v3

@@ II Non-commutative corrections for Schwarzschild black hole
Using the definition in Eq. (9), along with the above NC tetrad fields, and we stop up to the second order in the NC parameter $\Theta$, one can derive the following nonzero components of the deformed metric $\hat{g}_{\mu\nu}$ in the equatorial plane $\theta=\pi/2$:
$-\hat{g}_{tt}=$ $\left(1-\frac{2m}{r}\right)+\left\{\frac{m\left(88m^{2}+mr\left(-77+15\sqrt{1-\frac{2m}{r}}\right)-8r^{2}\left(-2+\sqrt{1-\frac{2m}{r}}\right)\right)}{16r^{4}(-2m+r)}\right\}\Theta^{2}+\mathcal{O}(\Theta^{4})$ (13a)
$\hat{g}_{rr}=$ $\left(1-\frac{2m}{r}\right)^{-1}+\left\{\frac{m\left(12m^{2}+mr\left(-14+\sqrt{1-\frac{2m}{r}}\right)-r^{2}\left(5+\sqrt{1-\frac{2m}{r}}\right)\right)}{8r^{2}(2m-r)^{3}}\right\}\Theta^{2}+\mathcal{O}(\Theta^{4})$ (13b)
$\hat{g}_{\theta\theta}=$ $r^{2}+\left\{\frac{m\left(m\left(10-6\sqrt{1-\frac{2m}{r}}\right)-\frac{8m^{2}}{r}+r\left(-3+5\sqrt{1-\frac{2m}{r}}\right)\right)}{16(-2m+r)^{2}}\right\}\Theta^{2}+\mathcal{O}(\Theta^{4})$ (13c)
$\hat{g}_{\phi\phi}=$ $r^{2}+\left\{\frac{5}{8}-\frac{3}{8}\sqrt{1-\frac{2m}{r}}+\frac{m\left(-17+\frac{5}{\sqrt{1-\frac{2m}{r}}}\right)}{16r}+\frac{m^{2}\sqrt{1-\frac{2m}{r}}}{(-2m+r)^{2}}\right\}\Theta^{2}+\mathcal{O}(\Theta^{4})$ (13d)
It is evident that, in the limit $\Theta\to 0$, the commutative Schwarzschild solution is recovered. A detailed discussion of the properties of the obtained metric can be found in Ref. abdellah1 for the equatorial plane, while the general case is discussed in Ref. abdellah2 .
@@
$\hat{g}_{00}=0\,,\quad 1/\hat{g}_{rr}=0.$ (14)
The solution to this equations give us respectively the NC static limit surface and the event horizon of the NC Schwarzschild BH abdellah1 ; abdellah2 :
$r_{sls}^{\mathrm{NC}}=r_{h}\left[1+\left(\frac{\Theta}{r_{h}}\right)\left(\frac{4\sqrt{5}+1}{32\sqrt{5}}\right)+\left(\frac{\Theta}{r_{h}}\right)^{2}\left(\frac{10+\sqrt{5}}{128}\right)\right],$ (15)
@@
$r_{h}^{NC}=r_{h}\left[1+\frac{3}{8}\left(\frac{\Theta}{r_{h}}\right)^{2}\right].$ (16)
It is clear that in the limit of $\Theta=0$, we recover the commutative event horizon $r_{h}=2m$.
III Geodetic motion and its stability around a non-commutative Schwarzschild black hole
In this section, we analyze the stability of circular orbits constrained to the equatorial plane (i.e., $\theta=\pi/2$) within the framework of the NC Schwarzschild spacetime. The corresponding spacetime metric is given by:
The motion of a test particle in the NC spacetime is governed by the Lagrangian:
@@ III Geodetic motion and its stability around a non-commutative Schwarzschild black hole
where the dot denotes differentiation with respect to the affine parameter $\tau$.
Applying the Euler-Lagrange equations, the generalized momenta are obtained as:
$p_{t}$ $=c^{2}\hat{g}_{tt}(r,\Theta)\dot{t}=E=\text{constant},$ (19a)
$p_{\phi}$ $=\hat{g}_{\phi\phi}(r,\Theta)\dot{\phi}=l=\text{constant},$ (19b)
$p_{r}$ $=\hat{g}_{rr}(r,\Theta)\dot{r}.$ (19c)
Since the Lagrangian is independent of $t$ and $\phi$, energy ($p_{t}=E$) and angular momentum ($p_{\phi}=l$) are conserved quantities. Consequently, the expressions for $\dot{t}$ and $\dot{\phi}$ become:
@@
Substituting these relations into the spacetime invariant equation $\hat{g}_{\mu\nu}U^{\mu}U^{\nu}\equiv-h$ (where $U^{\mu}=c^{-1}\frac{dx^{\mu}}{d\tau}$ is the four-velocity), we obtain the radial equation:
@@
where the effective potential $V_{\text{eff}}(r,\Theta)$ for a test particle in NC spacetime up to second order in the NC parameter $\mathbf{\Theta}$, is given by:
$V_{\text{eff}}(r,\Theta)$ $=\left(1-\frac{2m}{r}\right)\left(\frac{l^{2}}{r^{2}}+hc^{2}\right)-\frac{E^{2}}{c^{2}}+\Theta^{2}\left\{-\frac{l^{2}}{r^{4}}\left(1-\frac{2m}{r}\right)\left(\frac{5}{8}-\frac{3}{8}\sqrt{1-\frac{2m}{r}}+\frac{m\left(-17+\frac{5}{\sqrt{1-\frac{2m}{r}}}\right)}{16r}\right.\right.$
$\left.\left.+\frac{m^{2}\sqrt{1-\frac{2m}{r}}}{(-2m+r)^{2}}\right)+\frac{E^{2}}{c^{2}}\left(\frac{m(64m^{2}+m(-49+13\sqrt{1-\frac{2m}{r}})r+2(13-3\sqrt{1-\frac{2m}{r}})r^{2})}{16r^{5}(1-\frac{2m}{r})^{2}}\right)\right.$
$\left.+\left(\frac{l^{2}}{r^{2}}+hc^{2}\right)\left(\frac{m(12m^{2}+m(-14+\sqrt{1-\frac{2m}{r}})r-(5+\sqrt{1-\frac{2m}{r}}))r^{2}}{8r^{5}(1-\frac{2m}{r})}\right)\right\}+\mathcal{O}(\Theta^{4})$ (23)
In the commutative limit ($\Theta\to 0$), the effective potential reduces to the standard Schwarzschild form:
$V_{\text{eff}}(r,\Theta=0)=\left(1-\frac{2m}{r}\right)\left(\frac{l^{2}}{r^{2}}+hc^{2}\right)-\frac{E^{2}}{c^{2}}.$ (24)
@@ III.1 Radial motion of massive particles
$\left(\frac{dr}{d\tau}\right)^{2}$ $=E^{2}-\left(1-\frac{2m}{r}\right)-\Theta^{2}\left\{E^{2}\left(\frac{m(64m^{2}+m(-49+13\sqrt{1-\frac{2m}{r}})r+2(13-3\sqrt{1-\frac{2m}{r}})r^{2})}{16r^{5}(1-\frac{2m}{r})^{2}}\right)\right.$
$\quad\left.+\left(\frac{m(12m^{2}+m(-14+\sqrt{1-\frac{2m}{r}})r-(5+\sqrt{1-\frac{2m}{r}}))r^{2}}{8r^{5}(1-\frac{2m}{r})}\right)\right\}+\mathcal{O}(\Theta^{4}).$ (25)
@@
Next, we consider the free-fall motion of a massive particle in the NC Schwarzschild BH geometry. The particle is initially at rest, i.e., $\dot{r}=0$, at $r=r_{0}$ when $\tau=0$. For the proper time in NC spacetime, we set $E=1$ in Eq. (25) and integrate, yielding:
$\hat{\tau}$ $=-\int_{r_{0}}^{r}\left(\frac{2m}{r^{\prime}}-\Theta^{2}\left\{\left(\frac{m(64m^{2}+m(-49+13\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime}+2(13-3\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime 2})}{16r^{\prime 5}(1-\frac{2m}{r^{\prime}})^{2}}\right)\right.\right.$
$\quad\left.\left.+\left(\frac{m(12m^{2}+m(-14+\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime}-(5+\sqrt{1-\frac{2m}{r^{\prime}}}))r^{\prime 2}}{8r^{\prime 5}(1-\frac{2m}{r^{\prime}})}\right)\right\}\right)^{-1/2}dr^{\prime}.$
$=-\int_{r_{0}}^{r}\sqrt{\frac{2m}{r^{\prime}}}\left(1-\Theta^{2}\frac{r^{\prime}}{2m}\left\{\left(\frac{m(64m^{2}+m(-49+13\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime}+2(13-3\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime 2})}{16r^{\prime 5}(1-\frac{2m}{r^{\prime}})^{2}}\right)\right.\right.$
$\quad\left.\left.+\left(\frac{m(12m^{2}+m(-14+\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime}-(5+\sqrt{1-\frac{2m}{r^{\prime}}}))r^{\prime 2}}{8r^{\prime 5}(1-\frac{2m}{r^{\prime}})}\right)\right\}\right)^{-1/2}dr^{\prime}.$ (26)
@@
$\epsilon$ $=\Theta^{2}\frac{r^{\prime}}{2m}\left\{\left(\frac{m(64m^{2}+m(-49+13\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime}+2(13-3\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime 2})}{16r^{\prime 5}(1-\frac{2m}{r^{\prime}})^{2}}\right)\right.$
$\quad\left.+\left(\frac{m(12m^{2}+m(-14+\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime}-(5+\sqrt{1-\frac{2m}{r^{\prime}}}))r^{\prime 2}}{8r^{\prime 5}(1-\frac{2m}{r^{\prime}})}\right)\right\},$ (27)
where $\epsilon$ is a small parameter, due to the smallness of $\Theta$. Here we can use the first-order approximation of $\epsilon$:
@@
$\hat{\tau}$ $=\frac{2}{3}\left(\sqrt{\frac{r_{0}^{3}}{2m}}-\sqrt{\frac{r^{3}}{2m}}\right)+\Theta^{2}\left\{\sqrt{\frac{r_{0}}{2m}}\left(\frac{32m^{2}+4m(-52+\sqrt{1-\frac{2m}{r_{0}}})r_{0}+131r_{0}^{2}}{128(2m-r_{0})r_{0}^{2}}\right)\right.$
$\quad\left.-\sqrt{\frac{r}{2m}}\left(\frac{32m^{2}+4m(-52+\sqrt{1-\frac{2m}{r}})r+131r^{2}}{128(2m-r)r^{2}}\right)+\frac{1}{8m}\left(\text{ArcSin}\left(\sqrt{\frac{2m}{r_{0}}}\right)-\text{ArcSin}\left(\sqrt{\frac{2m}{r}}\right)\right)\right.$
$\quad\left.+\frac{67}{256m}\left(\text{ArcTanh}\left(\sqrt{\frac{r_{0}}{2m}}\right)-\text{ArcTanh}\left(\sqrt{\frac{r}{2m}}\right)\right)\right\}.$ (30)
@@
Fig. 1 illustrates the variation of proper time $\hat{\tau}$ as a function of $r$ during the free fall of a massive test particle into the NC Schwarzschild BH. The NC parameter $\Theta$ significantly influences the proper time near the event horizon, leading to an increase in its duration. The infinite time is required to reach the NC singularity resulting from non-commutativity effects (see Ref. abdellah1 ), and this phenomenon is absent in the commutative case Chandr1 .
Similarly to proper time, the coordinate time $\hat{t}$ for the free fall in the NC Schwarzschild BH can be obtained using (20) and (25) together with $E=1$:
$\hat{t}$ $=-\int_{r_{0}}^{r}\left(\hat{g}_{tt}\right)^{-1}\left(\frac{2m}{r^{\prime}}-\Theta^{2}\left\{\left(\frac{m(64m^{2}+m(-49+13\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime}+2(13-3\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime 2})}{16r^{\prime 5}(1-\frac{2m}{r^{\prime}})^{2}}\right)\right.\right.$
$\left.\left.+\left(\frac{m(12m^{2}+m(-14+\sqrt{1-\frac{2m}{r^{\prime}}})r^{\prime}-(5+\sqrt{1-\frac{2m}{r^{\prime}}}))r^{\prime 2}}{8r^{\prime 5}(1-\frac{2m}{r^{\prime}})}\right)\right\}\right)^{-1/2}dr^{\prime}\,,$ (31)
@@
$+\Theta^{2}\left\{\sqrt{\frac{r_{0}}{2m}}\left(-\frac{25}{16r_{0}}-\frac{37m}{128(r_{0}-2m)^{2}}+\frac{15\sqrt{1-\frac{2m}{r_{0}}}}{32(r_{0}-2m)^{2}}-\frac{r_{0}\sqrt{1-\frac{2m}{r_{0}}}}{4(r_{0}-2m)^{2}}-\frac{61}{512(r_{0}-2m)}\right)\right.$
$\left.-\sqrt{\frac{r}{2m}}\left(-\frac{25}{16r}-\frac{37m}{128(r-2m)^{2}}+\frac{15\sqrt{1-\frac{2m}{r}}}{32(r-2m)^{2}}-\frac{r\sqrt{1-\frac{2m}{r}}}{4(r-2m)^{2}}-\frac{61}{512(r-2m)}\right)\right.$
$\left.+\frac{1}{4m}\left(\text{ArcSin}\left(\sqrt{\frac{2m}{r_{0}}}\right)-\text{ArcSin}\left(\sqrt{\frac{2m}{r}}\right)\right)+\frac{605}{1024m}\left(\text{ArcTanh}\left(\sqrt{\frac{r_{0}}{2m}}\right)-\text{ArcTanh}\left(\sqrt{\frac{r}{2m}}\right)\right)\right\}.$ (32)
@@
In Fig. 2, we present the variation of coordinate time $\hat{t}$ during the free fall of a massive test particle in the NC Schwarzschild BH. The effect of non-commutativity become significant near the NC event horizon, where it lead to an increase in the coordinate time required for the particle to approach the NC horizon. This qualitative behavior aligns with the commutative case, as discussed in Chandr1 , but is further modified due to the presence of non-commutativity.
@@
Fig. 3 depicts the behavior of both proper time and coordinate time for a massive particle falling toward the NC Schwarzschild BH. These results indicate that non-commutativity prevents a massive particle from reaching the NC singularity within a finite time, in contrast to the commutative case Chandr1 . It is noteworthy that in this approach, the commutative case is recovered in the limit $\Theta=0$, and preserving the expected large-distance commutative behavior while introducing novel effects near the event horizon. This feature is not realized in other approaches rahaman ; bhar ; larranaga . Furthermore, the results highlight the impact of utilizing the gauge theory of gravity in NC spacetime to analyze the geodesic motion of test particles.
III.2 Circular motion of massive particles and Lyapunov exponents

### v4

@@ II Non-commutative corrections for Schwarzschild metric
Using the definition of Equation (10), together with the NC quadrupole fields, we obtain the non-zero components up to the second order of the NC parameter $\Theta$ for the deformed metric $\hat{g}_{\mu\nu}$ in the equatorial plane $\theta=\pi/2$. The NC line element of this new deformed spacetime can be written as follows:
where the NC metric component are given by $\hat{g}_{\mu\nu}$:
$-\hat{g}_{tt}=$ $\left(1-\frac{2m}{r}\right)+\left\{\frac{m\left(88m^{2}+mr\left(-77+15\sqrt{1-\frac{2m}{r}}\right)-8r^{2}\left(-2+\sqrt{1-\frac{2m}{r}}\right)\right)}{8r^{4}(-2m+r)}\right\}\Theta^{2}+\mathcal{O}(\Theta^{4})$ (14a)
$\hat{g}_{rr}=$ $\left(1-\frac{2m}{r}\right)^{-1}+\left\{\frac{m\left(-48m^{2}+55mr+4r^{2}\right)-4m^{2}r\sqrt{1-\frac{2m}{r}}+4mr^{2}\sqrt{1-\frac{2m}{r}}}{16r^{2}(r-2m)^{3}}\right\}\Theta^{2}+\mathcal{O}(\Theta^{4})$ (14b)
$\hat{g}_{\theta\theta}=$ $r^{2}+\left\{\frac{m\left(m\left(10-6\sqrt{1-\frac{2m}{r}}\right)-\frac{8m^{2}}{r}+r\left(-3+5\sqrt{1-\frac{2m}{r}}\right)\right)}{8(-2m+r)^{2}}\right\}\Theta^{2}+\mathcal{O}(\Theta^{4})$ (14c)
$\hat{g}_{\phi\phi}=$ $r^{2}+\left\{\frac{5}{8}-\frac{3}{8}\sqrt{1-\frac{2m}{r}}-\frac{2m}{r}+\frac{m\sqrt{1-\frac{2m}{r}}(-2m+7r)}{8(-2m+r)^{2}}\right\}\Theta^{2}+\mathcal{O}(\Theta^{4})$ (14d)
For a spherically symmetric Schwarzschild black hole in NC coordinates, the metric (14a-14d) exhibits anisotropy, which is physically reasonable. This anisotropy arises because noncommutativity effectively acts as a mass density within the BH, leading to an asymmetric distribution of this density along the coordinate axes. This asymmetry is a consequence of the deformed geometry of space, further reinforcing the anisotropic nature of the Schwarzschild metric in this context. Notably, two limiting cases can be observed: as $r$ approaches infinity, the metric (14a-14d) reduces to the Minkowski solution, and when the noncommutativity parameter $\Theta$ approaches zero, the metric (14a-14d) reverts to the standard Schwarzschild solution in the commutative case. Moreover, this geometry shifts the singularity at $r=0$ to a new finite radius, $r^{NC}_{\text{singularity}}=2m$ abdellah2 .
Interestingly, Eqs. (14a-14d) reveal that $\hat{g}_{tt}\hat{g}_{rr}\neq 1$, which is a necessary condition for satisfying all four energy conditions: the null energy condition, weak energy condition, dominant energy condition, and strong energy condition. This implies that the NC BH exhibits double photon spheres (for further details, see photonspheres ).
@@
$\hat{g}_{00}=0\,,\quad\frac{1}{\hat{g}_{rr}}=0.$ (15)
The solutions to these equations give us, respectively, the NC static limit surface and the event horizon of the NC Schwarzschild BH abdellah1 ; abdellah2 .
$r_{sls}^{\mathrm{NC}}=r_{h}\left[1+\left(\frac{\Theta}{r_{h}}\right)\left(\frac{4\sqrt{5}+1}{16\sqrt{10}}\right)+\left(\frac{\Theta}{r_{h}}\right)^{2}\left(\frac{10+\sqrt{5}}{64}\right)\right],$ (16)
@@
$r_{h}^{NC}=r_{h}\left[1+\frac{1}{4}\left(\frac{\Theta}{r_{h}}\right)^{2}\right].$ (17)
Our results differ from those reported in Ref. Tajron1 , due to the solution of the NC event horizon equation ($01/\hat{g}_{rr}=$ 0) up to the second order in $\Theta$, and different choices of tetrads fields. It is clear that in the limit of $\Theta=0$, we recover the commutative event horizon $r_{h}=2m$.
III Geodetic motion and stability of orbits around a non-commutative Schwarzschild black hole
In this section, we analyze the stability of circular orbits constrained to the equatorial plane (i.e., $\theta=\pi/2$) within the framework of the NC Schwarzschild spacetime. In this geometry (13), the motion of a test particle is governed by the following Lagrangian:
@@ III Geodetic motion and stability of orbits around a non-commutative Schwarzschild black hole
where the dot denotes differentiation with respect to the affine parameter $\lambda$.
Since the Lagrangian does not explicitly depend on $t$ and $\phi$, two conserved quantities arise along the path of the test particle, which correspond to the energy ($E$) and angular momentum ($l$) respectively. Consequently, the expressions for $\dot{t}$ and $\dot{\phi}$ are given by:
@@
Substituting these relations into the spacetime invariant equation:
$\hat{g}_{tt}U^{t}U^{t}+\hat{g}_{rr}U^{r}U^{r}+\hat{g}_{\phi\phi}U^{\phi}U^{\phi}\equiv-h$
where $U^{\mu}=c^{-1}\frac{dx^{\mu}}{d\tau}$ is the four-velocity, and with some algebra we obtain the radial equation:
@@
where the effective potential $V_{\text{eff}}(r,\Theta)$ for a test particle in NC spacetime up to second order in the NC parameter $\mathbf{\Theta}$, is given by abdellah1 :
$V_{\text{eff}}(r,\Theta)$ $=\left(1-\frac{2m}{r}\right)\left(\frac{l^{2}}{r^{2}}+hc^{2}\right)-\frac{E^{2}}{c^{2}}+\Theta^{2}\left\{-\frac{l^{2}}{r^{4}}\left(1-\frac{2m}{r}\right)\left(\frac{5}{8}-\frac{3}{8}\sqrt{1-\frac{2m}{r}}+\frac{m\sqrt{1-\frac{2m}{r}}(-2m+7r)}{8(-2m+r)^{2}}\right.\right.$
$\left.\left.-\frac{2m}{r}\right)+\frac{E^{2}}{c^{2}}\left(\frac{m\left(128m^{2}-12r^{2}\left(\sqrt{1-\frac{2m}{r}}-3\right)+mr\left(26\sqrt{1-\frac{2m}{r}}-99\right)\right)}{16r^{3}(r-2m)^{2}}\right)\right.$
$\left.+\left(\frac{l^{2}}{r^{2}}+hc^{2}\right)\left(\frac{m\left(48m^{2}-4r^{2}\left(\sqrt{1-\frac{2m}{r}}+1\right)+mr\left(4\sqrt{1-\frac{2m}{r}}-55\right)\right)}{16r^{4}(r-2m)}\right)\right\}+\mathcal{O}(\Theta^{4})$ (22)
In the commutative limit ($\Theta\to 0$), the effective potential reduces to the standard Schwarzschild form Chandr1 .
@@ III.1 Radial motion of massive particles
$\left(\frac{dr}{d\tau}\right)^{2}$ $=E^{2}-\left(1-\frac{2m}{r}\right)-\Theta^{2}\left\{E^{2}\left(\frac{m\left(128m^{2}-12r^{2}\left(\sqrt{1-\frac{2m}{r}}-3\right)+mr\left(26\sqrt{1-\frac{2m}{r}}-99\right)\right)}{16r^{3}(r-2m)^{2}}\right)\right.$
$\left.+\left(\frac{m\left(48m^{2}-4r^{2}\left(\sqrt{1-\frac{2m}{r}}+1\right)+mr\left(4\sqrt{1-\frac{2m}{r}}-55\right)\right)}{16r^{4}(r-2m)}\right)\right\}+\mathcal{O}(\Theta^{4}).$ (23)
@@
The effective potential in case of radial time-like geodesics is given by:
$V_{\text{eff}}(r,\Theta)$ $=\left(1-\frac{2m}{r}\right)-E^{2}+\Theta^{2}\left\{E^{2}\left(\frac{m\left(128m^{2}-12r^{2}\left(\sqrt{1-\frac{2m}{r}}-3\right)+mr\left(26\sqrt{1-\frac{2m}{r}}-99\right)\right)}{16r^{3}(r-2m)^{2}}\right)\right.$
$\left.+\left(\frac{m\left(48m^{2}-4r^{2}\left(\sqrt{1-\frac{2m}{r}}+1\right)+mr\left(4\sqrt{1-\frac{2m}{r}}-55\right)\right)}{16r^{4}(r-2m)}\right)\right\}+\mathcal{O}(\Theta^{4})$ (24)
Figure 1: Effective potential for a radial massive test particle as a function of $r$, for different values of $\Theta$ with fixed energy: $E=1$, $m=1$ (left panel), and for different energy $E$ with fixed $\Theta=0.2$ (right panel).
In Fig. 1, we illustrate the behavior of the effective potential for the radial trajectory ($l=0$) of a massive test particle in NC spacetime as a function of $r$. The left and right panels respectively depict variations for different values of the NC parameter and different energy levels of the particle. It is evident that, even in radial motion, non-commutativity prevents a massive test particle with high energy ($E=1$) from reaching the NC event horizon by predicting a new minimum outside the NC event horizon. Moreover, when we fix the NC parameter and vary the energy of the particle, an interesting behavior emerges: a particle with low energy can fall into the NC event horizon but cannot reach the NC singularity, which in this geometry is shifted to a finite radius, $r_{\text{singularity}}^{NC}=2m$. For a very low energy ($E\ll 1$), this phenomenon disappears, and the massive test particle falls into the NC singularity at the finite radius $2m$ (see Ref. abdellah1 ).
Next, we consider the free-fall motion of a massive particle in the NC Schwarzschild BH geometry. The particle is initially at rest, i.e., $\dot{r}=0$ at $r=r_{0}$ when $\tau=0$. For the proper time in NC spacetime, we set $E=1$ in Eq. (23) and integrate, yielding:
$\hat{\tau}$ $=-\int_{r_{0}}^{r}\left(\frac{2m}{r^{\prime}}-\Theta^{2}\left\{\left(\frac{m\left(128m^{2}-12r^{2}\left(\sqrt{1-\frac{2m}{r}}-3\right)+mr\left(26\sqrt{1-\frac{2m}{r}}-99\right)\right)}{16r^{3}(r-2m)^{2}}\right)\right.\right.$
$\quad\left.\left.+\left(\frac{m\left(48m^{2}-4r^{2}\left(\sqrt{1-\frac{2m}{r}}+1\right)+mr\left(4\sqrt{1-\frac{2m}{r}}-55\right)\right)}{16r^{4}(r-2m)}\right)\right\}\right)^{-1/2}dr^{\prime}.$
$=-\int_{r_{0}}^{r}\sqrt{\frac{2m}{r^{\prime}}}\left(1-\Theta^{2}\frac{r^{\prime}}{2m}\left\{\left(\frac{m\left(128m^{2}-12r^{2}\left(\sqrt{1-\frac{2m}{r}}-3\right)+mr\left(26\sqrt{1-\frac{2m}{r}}-99\right)\right)}{16r^{3}(r-2m)^{2}}\right)\right.\right.$
$\quad\left.\left.+\left(\frac{m\left(48m^{2}-4r^{2}\left(\sqrt{1-\frac{2m}{r}}+1\right)+mr\left(4\sqrt{1-\frac{2m}{r}}-55\right)\right)}{16r^{4}(r-2m)}\right)\right\}\right)^{-1/2}dr^{\prime}.$ (25)
@@
$\epsilon$ $=\Theta^{2}\frac{r^{\prime}}{2m}\left\{\left(\frac{m\left(128m^{2}-12r^{2}\left(\sqrt{1-\frac{2m}{r}}-3\right)+mr\left(26\sqrt{1-\frac{2m}{r}}-99\right)\right)}{16r^{3}(r-2m)^{2}}\right)\right.$
$\quad\left.+\left(\frac{m\left(48m^{2}-4r^{2}\left(\sqrt{1-\frac{2m}{r}}+1\right)+mr\left(4\sqrt{1-\frac{2m}{r}}-55\right)\right)}{16r^{4}(r-2m)}\right)\right\},$ (26)
where $\epsilon$ is a small parameter due to the smallness of $\Theta$. Here, we use the first-order approximation of $\epsilon$:
@@
$\hat{\tau}$ $=\frac{2}{3}\left(\sqrt{\frac{r_{0}^{3}}{2m}}-\sqrt{\frac{r^{3}}{2m}}\right)+\Theta^{2}\left\{\sqrt{\frac{r_{0}}{2m}}\left(\frac{64m^{2}+4m(-103+2\sqrt{1-\frac{2m}{r_{0}}})r_{0}+227r_{0}^{2}}{128(2m-r_{0})r_{0}^{2}}\right)\right.$
$\quad\left.-\sqrt{\frac{r}{2m}}\left(\frac{64m^{2}+4m(-103+2\sqrt{1-\frac{2m}{r}})r+227r^{2}}{128(2m-r)r^{2}}\right)+\frac{1}{4m}\left(\text{ArcSin}\left(\sqrt{\frac{2m}{r_{0}}}\right)-\text{ArcSin}\left(\sqrt{\frac{2m}{r}}\right)\right)\right.$
$\quad\left.+\frac{99}{256m}\left(\text{ArcTanh}\left(\sqrt{\frac{r_{0}}{2m}}\right)-\text{ArcTanh}\left(\sqrt{\frac{r}{2m}}\right)\right)\right\}.$ (29)
@@
Fig. 2 illustrates the variation of proper time $\hat{\tau}$ as a function of $r$ during the free fall of a massive test particle into the NC Schwarzschild BH. The NC parameter $\Theta$ significantly influences the proper time near the event horizon, increasing its duration. An infinite amount of time is required to reach the NC singularity due to non-commutativity effects (see Ref. abdellah1 ), a phenomenon that is absent in the commutative case Chandr1 .
Similarly to proper time, the coordinate time $\hat{t}$ for free fall in the NC Schwarzschild BH can be obtained using Eqs. (19) and (23) together with $E=1$:
$\hat{t}$ $=-\int_{r_{0}}^{r}\left(\hat{g}_{tt}\right)^{-1}\left(\frac{2m}{r^{\prime}}-\Theta^{2}\left\{\left(\frac{m\left(128m^{2}-12r^{2}\left(\sqrt{1-\frac{2m}{r}}-3\right)+mr\left(26\sqrt{1-\frac{2m}{r}}-99\right)\right)}{16r^{3}(r-2m)^{2}}\right)\right.\right.$
$\quad\left.\left.+\left(\frac{m\left(48m^{2}-4r^{2}\left(\sqrt{1-\frac{2m}{r}}+1\right)+mr\left(4\sqrt{1-\frac{2m}{r}}-55\right)\right)}{16r^{4}(r-2m)}\right)\right\}\right)^{-1/2}dr^{\prime}\,,$ (30)
@@
$+\Theta^{2}\left\{\sqrt{\frac{r}{2m}}\left(-\frac{3211m}{256(r-2m)^{2}}+\frac{25m^{2}}{2r(r-2m)^{2}}+\frac{1693r}{512(r-2m)^{2}}+\frac{(8r-15m)\sqrt{1-\frac{2m}{r}}}{16(r-2m)^{2}}\right)\right.$
$\left.-\sqrt{\frac{r_{0}}{2m}}\left(-\frac{3211m}{256(r_{0}-2m)^{2}}+\frac{25m^{2}}{2r_{0}(r_{0}-2m)^{2}}+\frac{1693r_{0}}{512(r_{0}-2m)^{2}}+\frac{(8r_{0}-15m)\sqrt{1-\frac{2m}{r_{0}}}}{16(r_{0}-2m)^{2}}\right)\right.$
$\left.+\frac{1}{2m}\left(\text{ArcSin}\left(\sqrt{\frac{2m}{r_{0}}}\right)-\text{ArcSin}\left(\sqrt{\frac{2m}{r}}\right)\right)+\frac{1181}{1024m}\left(\text{ArcTanh}\left(\sqrt{\frac{r_{0}}{2m}}\right)-\text{ArcTanh}\left(\sqrt{\frac{r}{2m}}\right)\right)\right\}.$ (31)
@@
In Fig. 3, we present the variation of coordinate time $\hat{t}$ during the free fall of a massive test particle in the NC Schwarzschild BH. The effect of non-commutativity becomes significant near the NC event horizon, where it leads to an increase in the coordinate time required for the particle to approach the NC horizon. This qualitative behavior aligns with the commutative case, as discussed in Chandr1 , but is further modified due to the presence of non-commutativity.
@@
Fig. 4 depicts the behavior of both proper time and coordinate time for a massive particle falling toward the NC Schwarzschild BH. These results indicate that non-commutativity prevents a massive particle from reaching the NC singularity within a finite time, in contrast to the commutative case Chandr1 , thus confirming our analysis of the effective potential of the radial trajectory of a massive test particle. It is noteworthy that in this approach, the commutative case is recovered in the limit $\Theta=0$, preserving the expected large-distance commutative behavior while introducing novel effects near the event horizon. This feature is not addressed in other approaches related to the NC matter distribution framework. In some cases, these approaches lose all commutative behavior (e.g., Refs. rahaman ; bhar ), while in others, they preserve the general commutative behavior but predict a reduction in proper time larranaga . This contrasts with our results, where this geometry ensures that the particle requires infinite proper time to reach the NC singularity. Furthermore, the results highlight the impact of utilizing the gauge theory of gravity in NC spacetime to analyze the geodesic motion of test particles.
III.2 Circular motion of massive particles